# An Efficient TLS 1.3 Handshake Protocol with VC Certificate Type

### v1

@@ V Implementation in OpenSSL
We have implemented the design presented in section III-B in OpenSSL, a globally adopted open source cryptographic library, to experimentally evaluate it. OpenSSL is mainly written in C language and consists of two sub-libraries: ssl that implements the TLS protocol and crypto that supplies a wide variety of cryptographic operations. Internally, crypto relies on loadable modules, called providers, that implement all the cyrptographic logic.
The main implementation goal is to minimize the changes to the OpenSSL sub-libraries to make smoother the contribution of the code upstream. In the end our implementation consists of slight modifications to the ssl module and no changes at all to the crypto library.
@@
In crypto, we treated the functions to deal with DIDs and VCs as cryptographic operations. We designed and implemented the ssi provider to supply these functions to OpenSSL and use them in the TLS handshake protocol. To leave the crypto library unchanged we mapped these functionalities into four existing operations:
@@
Fig. 3 shows the architecture of our overall solution. The implementation of the ssi provider operations relies on the identity-cbindings library. We developed the the identity-cbindings library as a wrapper of the IOTA Identity library [24]. The IOTA Identity library provides DID and VC-related functionalities in accordance to W3C specifications and the iota method to work with the IOTA Tangle  [25]. The purpose of the wrapper is to provide the ssi provider with C language interfaces to communicate with the IOTA Identity library, which is written in Rust language.
@@
The client and the server run the s_client and s_server applications respectively to start the TLS handshake. Both applications use the same cryptographic operations of the ssi provider. At the beginning these two applications load their identity and invoke the OSSL_OP_DECODING operation to decode them from PEM format and store them in IOTA Identity structures.
@@ VI-A Experimental Setup
To assess the performance of the new handshake we have installed the modified version of OpenSSL on two Raspberry Pi’s 4 (RPI) Model B equipped with a quad-core Cortex-A72 (ARM v8) SoC clocked at 1.8GHz, 4 GB of SDRAM, a Gigabit Ethernet interface, and 32-bit OS. The RPIs are connected in a client-server configuration and both have access to a DLT node to resolve the peer DID as depicted in Fig. 4. Without loosing generality, the RPIs leverage the immutability feature of the public IOTA Tangle [25] and interact with it through an IOTA node operated in our lab. The TLS endpoints resolve the peer DID with an HTTP request served by the IOTA node.
We tested the TLS handshakes between the two target RPIs under three different configurations for the connection between the RPIs and the IOTA node: (i) HTTP over TCP, (ii) HTTP over TCP with IPSec in endpoint-to-endpoint transport mode at layer 3, and (iii) HTTPs with a TLS channel established on the fly during the handshake between the RPIs. In the IPSec scenario the RPIs authenticate the IOTA node with an Ed25519 Raw Public Key, while in the TLS scenario the node employs a three-link X.509 certificate chain signed with a Ed25519 keys.
@@
In all tests, client and server adopt x25519 elliptic curve for Elliptic Curve Diffie-Hellman Ephemeral (ECDHE) key exchange and TLS_AES_256_GCM_SHA384 cipher suite [11]. We tested the performance of unilaterally authenticated handshake and mutually authenticated handshakes and compared them with the ones of the original handshake using a three-link X.509 certificate chain. We have only considered the Ed25519 signature algorithm as it is the algorithm supported by the IOTA Identity library [24]. It relies on the cryptographic implementation of Ed25519-zebra library [29]. We ran 1000 TLS handshakes for each configuration using s_client and s_server applications provided by OpenSSL to collect statistically relevant results about handshake size and latency. Client and server select randomly an X.509 certificate chain or a VC from a predefined large set at each run.
@@ VI-B Experimental results
Table I shows the number of bytes sent by the server to the client in a unilaterally authenticated handshake while using X.509 and VC certificates. The content, and so the number of bytes, in a VC or X.509 certificate can vary based on how much information they contain. Therefore, the total bytes column should be considered as indicative. Instead, the values about public key objects are fixed. In an handshake with VC certificates type a server sends half of the bytes of public key objects than an original handshake does. Note that in an unilaterally authenticated handshake with the VC certificate type, only the client resolves the server DID. Conversely, in a mutually authenticated handshake also the server interacts with the DLT node. In the latter scenario, when the server establish a TLS channel with the DLT node to resolve the client DID, it sends an additional 377 bytes. In general, endpoints may find more suitable IPSec alternative as the secure channel gets established before the handshake and so less bytes will be exchanged over the network. In the case of mutually authenticated handshakes the values about public key objects in Table I double.
@@
the upper part of Table II shows the average latency of an unilaterally authenticated handshake using either an X.509 or a VC certificate type.
@@
Overall, the experimental results show that the IPSec approach offers the best trade-off between security and performance. However, the performances in Table II represents the worst case. In fact, in an IoT system we suppose there is only one Issuer and its DID and public key can be stored on each IoT node halving the number of required DID Resolves, hence the overall latency of an handshake between the target IoT nodes. The DID Resolve operation would drop from 9,2 ms to 5,2 ms and the TLS handshake latency from 26,4 ms to 22,4 ms on average.
@@
Table III shows the average latency of hybrid TLS handshakes between the target IoT nodes. We tested the TLS handshakes while protecting the communication with the DLT node by means of IPSec. As expected the average latency are almost equal in both cases, since the sequence of operations performed are the same. In addition, we tested the fallback to the original handshake to cover the case a server does not implement RFC-7250 [16] or does not recognise the new VC certificate type. We installed two different versions of OpenSSL on the two IoT nodes; one with our implementation on the client and an original version on the server. The functional tests were successful. The average handshake latency remains the same as in the original handshake case.
@@ VII Conclusion and Future Work
To support and facilitate the adoption of SSI in large-scale IoT systems, we have taken a step forward in the design and implementation of a TLS 1.3 handshake protocol that enables the use of VCs while maintaining full compliance with RFC-8446 and preserving all the security features of TLS 1.3. The improvement over our previous work lies in the handshake design, which now only uses messages already defined for TLS 1.3 and a new extension. The design has an incredibly positive impact on the implementation, as we made minimal changes to the OpenSSL library and relied mostly on our external ssi provider to handle VC and DID-related operations. Furthermore, the experimental results prove the feasibility of our design and show comparable performance to the original solution based on PKI and X.509 certificates. These results pave the way for the adoption of SSI in large-scale IoT systems, with a clear benefit in terms of reducing the cost of identity management. Future work will focus on further optimising the implementation to further improve performance.

### v2

@@ V Implementation in OpenSSL
We have implemented the design presented in section III-B in OpenSSL, a globally adopted open source cryptographic library, to experimentally evaluate it. OpenSSL is mainly written in C language and consists of two sub-libraries: ssl that implements the TLS protocol and crypto that supplies a wide variety of cryptographic operations. Internally, crypto relies on loadable modules, called providers, that implement all the cryptographic logic.
The main implementation goal is to minimize the changes to the OpenSSL sub-libraries to make smoother the contribution of the code upstream. In the end, our implementation consists of slight modifications to the ssl module and no changes at all to the crypto library.
@@
In crypto, we treated the functions to deal with DIDs and VCs as cryptographic operations. We designed and implemented the ssi provider to supply these functions to OpenSSL and use them in the TLS handshake protocol. To leave the crypto library unchanged, we mapped these functionalities into four existing operations:
@@
Fig. 3 shows the architecture of our overall solution. The implementation of the ssi provider operations relies on the identity-cbindings library. We developed the identity-cbindings library as a wrapper of the IOTA Identity library [24]. The IOTA Identity library provides DID and VC-related functionalities in accordance to W3C specifications and the iota method to work with the IOTA Tangle  [25]. The purpose of the wrapper is to provide the ssi provider with C language interfaces to communicate with the IOTA Identity library, which is written in Rust language.
@@
The client and the server run the s_client and s_server applications respectively to start the TLS handshake. Both applications use the same cryptographic operations of the ssi provider. At the beginning, these two applications load their identity and invoke the OSSL_OP_DECODING operation to decode them from PEM format and store them in IOTA Identity structures.
@@ VI-A Experimental Setup
To assess the performance of the new handshake we have installed the modified version of OpenSSL on two Raspberry Pi’s 4 (RPI) Model B equipped with a quad-core Cortex-A72 (ARM v8) SoC clocked at 1.8GHz, 4 GB of SDRAM, a Gigabit Ethernet interface, and 32-bit OS. The RPIs are connected in a client-server configuration and both have access to a DLT node to resolve the peer DID as depicted in Fig. 4. Without loosing generality, the RPIs leverage the immutability feature of the public IOTA Tangle [25] and interact with it through an IOTA node operated in our lab. The TLS endpoints resolve the peer DID with an HyperText Transfer Protocol (HTTP) request served by the IOTA node.
We tested the TLS handshakes between the two target RPIs under three different configurations for the connection between the RPIs and the IOTA node: (i) HTTP over TCP, (ii) HTTP over TCP with IPSec in endpoint-to-endpoint transport mode at layer 3, and (iii) HTTPs with a TLS channel established on the fly during the handshake between the RPIs. In the IPSec scenario, the RPIs authenticate the IOTA node with an Ed25519 Raw Public Key, while in the TLS scenario the node employs a three-link X.509 certificate chain signed with Ed25519 keys.
@@
In all tests, client and server adopt x25519 elliptic curve for Elliptic Curve Diffie-Hellman Ephemeral (ECDHE) key exchange and TLS_AES_256_GCM_SHA384 cipher suite [11]. We tested the performance of unilaterally authenticated handshake and mutually authenticated handshakes and compared them with the ones of the original handshake using a three-link X.509 certificate chain. We have only considered the Ed25519 signature algorithm, as it is the algorithm supported by the IOTA Identity library [24]. It relies on the cryptographic implementation of Ed25519-zebra library [29]. We ran 1000 TLS handshakes for each configuration using s_client and s_server applications provided by OpenSSL to collect statistically relevant results about handshake size and latency. Client and server select randomly an X.509 certificate chain or a VC from a predefined large set at each run.
@@ VI-B Experimental results
Table I shows the number of bytes sent by the server to the client in a unilaterally authenticated handshake while using X.509 and VC certificates. The content, and so the number of bytes, in a VC or X.509 certificate can vary based on how much information they contain. Therefore, the total bytes column should be considered as indicative. Instead, the values about public key objects are fixed. In a handshake with VC certificate type, a server sends half of the bytes of public key objects than an original handshake does. Note that in a unilaterally authenticated handshake with the VC certificate type, only the client resolves the server DID. Conversely, in a mutually authenticated handshake also the server interacts with the DLT node. In the latter scenario, when the server establishes a TLS channel with the DLT node to resolve the client DID, it sends an additional 377 bytes. In general, endpoints may find more suitable the IPSec alternative as the secure channel gets established before the handshake and so fewer bytes will be exchanged over the network. In the case of mutually authenticated handshakes, the values about public key objects in Table I double.
@@
the upper part of Table II shows the average latency of a unilaterally authenticated handshake using either an X.509 or a VC certificate type.
@@
Overall, the experimental results show that the IPSec approach offers the best trade-off between security and performance. However, the performances in Table II represents the worst case. In fact, in an IoT system we suppose there is only one Issuer and its DID and public key can be stored on each IoT node halving the number of required DID Resolves, hence the overall latency of a handshake between the target IoT nodes. The DID Resolve operation would drop from 9,2 ms to 5,2 ms and the TLS handshake latency from 26,4 ms to 22,4 ms on average.
@@
Table III shows the average latency of hybrid TLS handshakes between the target IoT nodes. We tested the TLS handshakes while protecting the communication with the DLT node by means of IPSec. As expected, the average latency are almost equal in both cases, since the sequence of operations performed are the same. In addition, we tested the fallback to the original handshake to cover the case a server does not implement RFC-7250 [16] or does not recognize the new VC certificate type. We installed two different versions of OpenSSL on the two IoT nodes; one with our implementation on the client and an original version on the server. The functional tests were successful. The average handshake latency remains the same as in the original handshake case.
@@ VII Conclusion and Future Work
To support and facilitate the adoption of SSI in large-scale IoT systems, we have taken a step forward in the design and implementation of a TLS 1.3 handshake protocol that enables the use of VCs while maintaining full compliance with RFC-8446 and preserving all the security features of TLS 1.3. The improvement over our previous work lies in the handshake design, which now only uses messages already defined for TLS 1.3 and a new extension. The design has an incredibly positive impact on the implementation, as we made minimal changes to the OpenSSL library and relied mostly on our external ssi provider to handle VC and DID-related operations. Furthermore, the experimental results prove the feasibility of our design and show comparable performance to the original solution based on PKI and X.509 certificates. These results pave the way for the adoption of SSI in large-scale IoT systems, with a clear benefit in terms of reducing the cost of identity management. Future work will focus on further optimizing the implementation to further improve performance.
Acknowledgment
This work has been developed within the SEDIMARK project (https://sedimark.eu), funded by the European Union under the Horizon Europe framework programme [grant agreement no. 101070074].